OpenStudy (anonymous):

What is the inverse of the equation y = 2x - 3?

OpenStudy (zzr0ck3r):

add 3 to both sides then divide both sides by 2 then tell me what you have

OpenStudy (abb0t):

solve for "x" in terms of "y".

OpenStudy (anonymous):

you switch the x and the y

OpenStudy (anonymous):

x<2 ?

OpenStudy (zzr0ck3r):

I think you need to go back a few sections.

OpenStudy (anonymous):

y = 2x - 3 x=2y -3 y= (-x -3)/-2 y= x/2 + 3/2

OpenStudy (zzr0ck3r):

this is not going to help someone who thought the answer was x<2 at this point you are just doing their homework for them @hardequalsmath

OpenStudy (anonymous):

i showed him how to do it, thats what I would of wanted so now he could know how to do the rest